OpenStudy (mendicant_bias):

(Orthogonal Functions) More info below, finding the orthogonality of a complex function in the Hermitian sense.

OpenStudy (mendicant_bias):

\[\text{Show that the set $e^{inx}, -\pi \leq x \leq \pi, n \in Z,$ is orthogonal} \]

OpenStudy (mendicant_bias):

(Orthogonal in the Hermitian sense)

OpenStudy (mendicant_bias):

My idea of this would be that if \[w(x)=u(x)+iv(x)\]is the general form of a complex function, then in this situation \[iv(x)=e^{inx}\]\[u(x)=e^0\] (Need to find the complex conjugate to compute the Hermitian) That just sounds fishy to me, though. Would that really just mean that its complex conjugate is 1? @SithsAndGiggles

OpenStudy (mendicant_bias):

The way I'm currently imagining this is that\[u(x)=0, \ \ \ iv(x)=inx, \ \ \ \overline{w}(x)=u(x)-iv(x)=e^{0-inx}=e^{-inx}\]

OpenStudy (mendicant_bias):

Is this right? @perl

OpenStudy (mendicant_bias):

@Zarkon

OpenStudy (mendicant_bias):

(Misread originally the definition of a complex function, put things as I understand them with the correct definition in the fourth from current post)

OpenStudy (perl):

demonstrate that $$ \Large { \int_{-\pi}^{\pi}e^{inx}\cdot \overline { e^{inx} }=0 } $$

OpenStudy (mendicant_bias):

I'm not even sure about the definition; what I want to understand before I do anything else is whether the complex conjugate of \[e^{inx}\] is\[e^{-inx}\]

OpenStudy (mendicant_bias):

Also, how did you scale up the font? I'm not sure how to do that in the OS LaTeX editor.

OpenStudy (mendicant_bias):

Alright, so that answers my question, cool. Then you just have the definite integral of zero, which is c, but since you're evaluating at those bounds, it's just c-c=0.

OpenStudy (perl):

oh wait, its not so easy

OpenStudy (mendicant_bias):

(?)

OpenStudy (mendicant_bias):

But yeah, how did you scale up the font in the OS LaTeX editor, I need to know how to do that.

OpenStudy (perl):

type \Large , or \large

OpenStudy (perl):

you can use brackets around multiple sentences \Large { formulas and stuff }

OpenStudy (mendicant_bias):

Alright, awesome, I wasn't aware that would work in this environment. But yeah, why is it not so easy/what is wrong with what I said?

OpenStudy (perl):

$$ \Large { \Large { \int_{-\pi}^{\pi}e^{inx}\cdot \overline { e^{imx} }=0 \\\text { } \\ \text{however}\\\text { } \\ \overline { e^{imx} } = e^{-imx} } } $$

OpenStudy (perl):

for any m not equal to n

OpenStudy (perl):

scroll up one page to get the definition of Hermitian orthogonal

OpenStudy (mendicant_bias):

I get an HTTP 500 error, but I'll try again.

OpenStudy (perl):

one sec, i see

OpenStudy (mendicant_bias):

Yeah, no, I agree with you on that, heh, I just am not sure how to do it from here.

OpenStudy (perl):

then scroll up a page to get the definition of hermitian

OpenStudy (mendicant_bias):

Nah, it worked the second time for whatever reason, and I'm in agreement/knew that, just thinking about how to now demonstrate that with m =/= n.

OpenStudy (perl):

we have to solve that integral

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}e^{inx}e^{-imx}=e^{n/m}dx\]

OpenStudy (mendicant_bias):

\[\large{ \int\limits_{-\pi}^{\pi}e^{n/m}dx=xe^{n/m}\bigg]^{\pi}_{-\pi} }\]

OpenStudy (mendicant_bias):

Ah, nevermind, I see what you're doing, my bad

OpenStudy (perl):

$$ \Large { \int_{-\pi}^{\pi}e^{inx}\cdot \overline { e^{imx}} ~dx \\= \int_{-\pi}^{\pi}e^{inx}\cdot { e^{-imx} }~dx\\ =\int_{-\pi}^{\pi}e^{inx-imx}~dx\\ =\int_{-\pi}^{\pi}e^{i(n-m)x}~dx\\ } $$

OpenStudy (mendicant_bias):

\[\large{\int\limits_{-\pi}^{\pi}e^{(ix)(n-m)}dx }\]

OpenStudy (perl):

if you write click on my latex code, you can see 'show text as ' and choose TEx

OpenStudy (mendicant_bias):

(Lol for whatever reason didn't read your last line, will try to integrate/move forward from here) Alright

OpenStudy (perl):

choose "show math as" choose Tex command. then you can copy paste (begin and end the latex code with double dollar signs)

OpenStudy (perl):

i never use the equation editor. all the latex commands are online anyway

OpenStudy (perl):

ok now we integrate this

OpenStudy (mendicant_bias):

\[\Large{\int\limits_{-\pi}^{\pi}e^{(m-n)ix}dx= \left[ \frac{e^{(m-n)ix}}{(m-n)i} \right] }\]

OpenStudy (mendicant_bias):

I'm not sure why, as opposed to using TeX Studio, where that doesn't happen, the left bracket isn't disappearing when I go "\left.", but in any case

OpenStudy (perl):

$$ \Large { \int_{-\pi}^{\pi}e^{inx}\cdot \overline { e^{imx}} ~dx \\= \int_{-\pi}^{\pi}e^{inx}\cdot { e^{-imx} }~dx\\ =\int_{-\pi}^{\pi}e^{inx-imx}~dx\\ =\int_{-\pi}^{\pi}e^{i(n-m)x}~dx\\ u = i(n-m)x\\ du = i(n-m)~dx\\ \frac{du}{i(n-m)}= dx \\ \int_{-\pi}^{\pi}e^{i(n-m)x}~dx = \left [\frac{e^{(n-m)ix} } {(n-m)i} \right]_{-\pi}^{\pi} } $$

OpenStudy (perl):

i tried the same, to delete the left bracket. wont allow it

OpenStudy (mendicant_bias):

Sorry, had to go to class. But alright, looks like this makes sense.